OpenStudy (anonymous):

The manufacturing cost of a cell phone was $136. Sellers marked up the cost by 25%. What was the new selling price? A-$170 B-$150.12

OpenStudy (anonymous):

136/4=34 136+34=170

OpenStudy (anonymous):

Or... 136*0.25=34 136+34=170

OpenStudy (anonymous):

.25x136=34 34+136=170 Answer is A

OpenStudy (anonymous):

Or 136(1+.25) = answer